# Verifiable blind observable estimation: A composably secure protocol for near-term quantum advantage tasks

### v1

@@ Abstract
The rapid advance of quantum hardware is spotlighting pre-fault-tolerant tasks that may no longer be efficiently validated by classical means and are likely to run on potentially untrusted remote quantum servers. This motivates problem-independent verification protocols with rigorous guarantees. The Verifiable Blind Quantum Computation (VBQC) protocol provides delegated computation where the composable security spans the confidentiality and integrity of the computation.
However, the success of these cryptographic protocols, especially their low space overhead, is unfortunately confined to problems that admit an algorithm whose output can be amplified through majority voting toward the correct solution. This leaves various notable near-term applications relying on observable estimation without efficient verification protocols.
To address these needs, we introduce a protocol implementing Secure Delegated Observable Estimation (SDOE), which efficiently verifies observable estimation performed on an untrusted quantum machine. More precisely, it guarantees that the computed estimate is within some $\epsilon>0$ of the true expectation value or else it aborts. The required overhead is limited to adding test rounds that are not more complex than the unprotected computation that needs to be performed to implement the desired measurement on a given fiducial state; and in addition, the security error is negligible in the total number of rounds of the protocol.
@@ I Introduction
Quantum computing has long been anticipated as a new computational paradigm capable of solving problems that are considered intractable for classical computation, referred to as quantum advantage [1, 2, 3, 4, 5, 6]. With the rapid advancement of quantum hardware [7, 8], increasing attention has been directed towards tasks that may demonstrate quantum advantage without requiring full fault tolerance [9, 10, 11, 12]. These are, by construction, quantum tasks that produce outputs that are classically intractable to verify. The difficulty is that the limitations and imperfections of near-term quantum hardware make it difficult to ensure reliable computation even if an algorithm has a provable quantum advantage. In addition, the recent increase in remote access to quantum devices poses the question of privacy and integrity of delegated computation. These considerations indicate the need for methods that enable efficient verification and secure delegation of quantum computations onto quantum devices beyond the direct control of an end user.
The verifiable blind quantum computation (VBQC) protocols [13, 14, 15] address this task theoretically and achieve unconditional verifiability and perfect blindness using measurement-based quantum computation (MBQC) [16, 17, 18, 19, 20], when the client can prepare and transmit single qubits to the server. Using trap qubits to verify the computation, the soundness error of the VBQC protocol [13] can be made exponentially small, while it requires full fault-tolerance to exponentially amplify its security. Recent work introducing the Robust VBQC (RVBQC) protocol [14] that verifies BQP computations further removes the qubit overhead for trap embedding and security amplification, making the protocols more practical [21, 22]. All these protocols are composably secure within the framework of abstract cryptography (AC) [23, 24, 25], allowing them to be integrated in the full-stack software combined with other protocols one wishes to run.
However, the success of these cryptographic protocols has been confined to problems that admit an algorithm whose output can be amplified through majority voting toward the correct solution. This reveals a fundamental gap, as the majority of algorithms driving the quest for quantum advantage, from quantum simulation [26, 27] to machine learning [27, 28, 29, 30], rely on expectation value estimation, which does not possess this structure on a per-round basis. Applying existing verification techniques to these tasks then becomes impractical, typically requiring inefficient methods like binary search with a sequence of decision tasks that add significant circuit overhead [31]. This has left the most common quantum applications without a path to achieving composable security with efficient verification and negligible soundness error.
In this work, we propose the Verifiable Blind Observable Estimation (VBOE) protocol to address the client’s need for obtaining and verifying the expectation value of an observable with a bias lower than a tolerable threshold defined by the client, in a way that preserves the confidentiality of the measured observable and returned estimation. To this end, we introduce a new ideal resource, which we call the Secure Delegated Observable Estimation (SDOE) resource, that captures the desired functionality. Our main result then consists of a proof that VBOE constructs SDOE with negligible error in the number of rounds that are used within the protocol.
The overall structure of VBOE follows that of RVBQC, randomly interleaving computation rounds and test rounds, where the Client either accepts or aborts the computation based on the number of failed test rounds. Our protocol, however, differs from RVBQC in its classical post-processing: by directly averaging the outputs of computation rounds, VBOE estimates the expectation value without the circuit overhead required by previous approaches.
Figure 1: A summary of our contribution. The upper part depicts the mutually beneficial relationship between verification protocols, observable estimation tasks, and error mitigation methods towards secure verifiable quantum advantage. The lower part sketches the exponential security of VBOE within the AC framework characterised by the distinguishability of the protocol and our new ideal resource (SDOE) capturing the desired verification of observable estimation tasks.
@@ II Verifying observable estimation
Despite the advantages of RVBQC in reducing overhead and accommodating noise-robustness, its effective applicability remains limited to problems that admit an algorithm whose result can be amplified through a majority vote. In practice, many relevant quantum tasks, such as quantum simulation, variational algorithms, and learning problems, involve estimating expectation values of observables with a given bias. While this can be made through RVBQC, this would require computing the empirical average over measurement shots in a quantum fashion. Here, instead, we define the Secure Delegated Observable Estimation (SDOE) resource (Resource 1) and show that it can be constructed by a protocol (Protocol 1) to negligible error within the Abstract Cryptography framework.
@@ II.1 Observable estimation problems
Using Hoeffding’s bound, one can assess the performance of such a procedure:
[…]
This motivates the following definition:
@@
[…]
@@ II.2 Secure delegated observable estimation (SDOE)
[…]
Resource 1 (Secure Delegated Observable Estimation (SDOE)).
Public information: $\mathfrak{C}$ a computation class; $N_{c},N_{t}\in\mathbb{N}$ and $w>0$ some security parameters; and $\epsilon>0$ the allowed bias.
Client’s interface: The target computation $\mathsf{C}\in\mathfrak{C}$ to produce the single qubit state $\rho$ to be measured by $|O=\outerproduct{-}{-}$.
@@
    If $e=1$, it computes $s$ using the transmitted state $\sigma$ and $F$
@@
    If $s=\mathsf{Abort}$ it forwards $o=\mathsf{Abort}$ to the Client
@@
[…]
@@
The pictorial representation of this resource is provided in Fig. 2.
[…]
This definition corresponds to the intuitive notion of secure delegated observable estimation: a malicious Server can only learn the class of observable estimation problems that the resource can tackle, and is only able to influence the result so long as it remains within $\epsilon$ of the true expectation value. It also inevitably has the ability to force the protocol to $\mathsf{Abort}$.
@@ II.3 Verifiable blind observable estimation (VBOE)
Considering the SDOE resource defined above, we propose the VBOE protocol to concretely implement this functionality. It is based on the same sequential execution of test and computation rounds as in RVBQC, with post-processing and acceptance criteria tailored to observable estimation problems. Here we assume that $\mathfrak{C}$ is a computation class that corresponds to all measurement patterns that can be executed on a graph $G=(V,E)$ with a given flow $f$.
Protocol 1 (Verifiable Blind Observable Estimation (VBOE)).
[…]
@@
    For $i\in\mathrm{S}_{\mathsf{T}}$, the Client constructs a test round following the same procedure as for the test rounds of the RVBQC (Protocol 5).
@@
    Each round, computation or test, is then delegated to the Server using UBQC (Protocol 3)
@@
    Upon receiving and decoding the result of the computation round $i\in\mathrm{S}_{\mathsf{C}}$, the Client assigns the result to $\tilde{y}_{i}$
@@
    If less than $wN_{t}$ test rounds failed, it sets $\tilde{o}=\frac{1}{N_{c}}\sum_{i\in S_{\mathsf{C}}}\tilde{y}_{i}$ as the result, otherwise it sets it to $\tilde{o}=\mathsf{Abort}$.
The schematic illustration of the VBOE protocol is depicted in Fig. 3.
@@
The main modification in VBOE from RVBQC is the post-processing of computation rounds. RVBQC applies classical majority vote over multiple repeated runs to amplify the probability of choosing a correct outcome, while VBOE computes the empirical average over $N_{c}$ outcomes $\{\tilde{y}_{i}\}_{i\in S_{\mathsf{C}}}$. To ensure that the returned value stays within the allowed bound $\epsilon$, the threshold $w$ needs to be set at the appropriate value. More precisely, we will see in the security proof that it is set in a way that the concentration of probabilities ensures honest executions are always accepted, while deviations that could generate a result too far away from the true value are rejected.
@@ II.4 Concrete construction of the SDOE resource
Let $\mathfrak{C}$ be a class of observable estimation problems that can be estimated using an MBQC pattern on a fixed graph $G$ with a given flow $f$ and chromatic number $k$. Let $N_{c},N_{t},\in\mathbb{N}$, let $\epsilon,w$ be constants such that $0\leq kw<\epsilon$. Then, the VBOE protocol (Protocol 1), with $N_{c}$ computation rounds and $N_{t}$ test rounds, $\delta$-constructs the SDOE resource. For a constant ratio $N_{c}/N_{t}$, $\delta$ is negligible in $N_{c}$.
Following abstract cryptography, to prove this theorem, we need to upper bound the distinguishing advantage between the VBOE protocol and the SDOE resource in the honest (Correctness proof) and malicious (Security proof) settings (see Definition 3).
@@
The proof of correctness relies on the composability of the UBQC protocol. As apparent in Protocol 1, each round is delegated to the server using UBQC (Protocol 3). Because UBQC perfectly constructs the Blind Delegated Quantum Computation (Resource 2), we can instead perform the proof of correctness using a hybrid protocol where each instantiation of UBQC is replaced by a call to BDQC.
[…]
We can thus conclude that, for $\epsilon$ fixed, the distinguishing advantage in the honest case is a negligible function of $N_{c}$. ∎
@@
As for the correctness proof, security relies heavily on the composability of UBQC.
We start by constructing a simulator that we attach to the Server’s interface of the ideal resource. Its purpose is to generate plausible transcripts and help the ideal resource in returning $o$ to the Client’s interface so that a distinguisher will be unable to tell apart this situation from running the Client’s part of the concrete Protocol 1. This simulator is easily constructed from the simulator designed to prove the security of UBQC.
First, it sets $e=1$. It then prepares EPR pairs for each qubit that the Server is supposed to receive in Protocol 1. It sends all half EPR-pairs to the Server, instructs random measurement angles, and retrieves alleged measurement outcomes. It then forwards the second half of the EPR pairs, the chosen angles and received bits to the ideal resource. It also samples at random indices within $N_{c}+N_{t}]$ to define the sets $\mathrm{S}_{\mathsf{C}}$ and $\mathrm{S}_{\mathsf{T}}$, it decides which type of test round is associated with each $i\in\mathrm{S}_{\mathsf{T}}$ and passes this to the ideal resource. Following the security proof of UBQC, the information passed per round, together with the nature of the round—computation or test—is sufficient for the ideal resource to generate per-round measurement results following the same, possibly deviated, probability distributions as the ones obtained when running Protocol 1. From the outcomes of computation rounds, the ideal resource is then instructed to compute the empirical average $o$ and from the test rounds to check that no more than $wN_{t}$ failed, in which case it sets $o=\mathsf{Abort}$.
[…]
As a result, the distinguishing advantage for telling apart the concrete protocol from the ideal resource is bounded by the total variation distance between the probability distributions of $s$ and $o$, or equivalently between $\tilde{o}$ and $o$. Because, conditioned on $\tilde{o},o\in\mathbb{R}$ or $\tilde{o},o=\mathsf{Abort}$ the distributions of $\tilde{o}$ and $o$ are the same, the total variation distance reduces to:
[…]
@@
[…]
@@
    given that $\tilde{o}\neq\mathsf{Abort}$, we then upper-bound the probability that a large number of computation rounds, say $(kw+\gamma_{2})N_{c}$ for $\gamma_{2}>0$, have been attacked by the server;
@@
    we recognise that attacking a fraction $\phi$ of computation rounds yields a deviated empirical average that is not away by more than $\phi$ from the non-deviated one, due to the binary nature of the observable $O$;
@@
[…]
@@
The probability at step 1 is upper-bounded by $2\exp(-2\gamma_{1}^{2}N_{c})$ using Hoeffding’s inequality.
@@
The probability of step 2 is upper-bounded in the following way. First, because UBQC reduces the attack by the Server to a convex combination of Pauli deviations before the measurements, we can group the deviation strategies by $m$, the number of attacked rounds. Now, because the location of test rounds and computation rounds are random, the number of affected computation rounds $Z$ follows a hypergeometric distribution with parameters $N_{c}+N_{t}$ for the total number of items, $m$ for the number of marked items and $N_{c}$ for the number of samples. Similarly, the number of affected test rounds $X$ follows a hypergeometric distribution with the roles $N_{c}$ and $N_{t}$ swapped.
@@
The idea is now to upper-bound the probability $\Pr[Z\geq(kw+\gamma_{2})N_{c},Y\leq w]$ using the fact that, for $m_{0}=(kw+\gamma_{2}/2)(N_{c}+N_{t})$, we have

### v2

@@ Abstract
Cryptographic verification is essential for establishing trust in quantum-computing-as-a-service. However, a fundamental gap exists in the current verification landscape: existing efficient protocols are largely restricted to decision problems where correctness is boosted by classical majority voting. This excludes observable estimation, the statistical task underpinning nearly all near-term quantum advantage applications. For such tasks, current verification techniques face a prohibitive trade-off: either weak security guarantees or massive space overhead that exceeds the capacity of near-term hardware. To resolve this, we introduce the Secure Delegated Observable Estimation (SDOE) ideal resource, the first formal cryptographic framework for trustworthy expectation-value estimation within Abstract Cryptography. We then present the Verifiable Blind Observable Estimation (VBOE) protocol, which efficiently constructs this resource. VBOE circumvents the limitations inherent in prior methodologies by enabling the sequential collection of samples with negligible security error, requiring zero extra qubit overhead. By directly averaging computation rounds in classical post-processing, our protocol provides the only known path to rigorous, composable verification for the most common class of near-term quantum-classical hybrid algorithms. This work bridges foundational cryptographic theory with practical quantum tasks, enabling the certification of quantum utility on current and near-future devices.
@@ I Introduction
Quantum computing promises to solve problems intractable for classical machines, a capability known as quantum advantage [Shor1994Algorithms, Grover1996A, Preskill2018quantumcomputingin, Preskill2012Quantum, Lanes2025A, Huang2025The]. Recent advances in quantum hardware [ni2023beating, gupta2024encoding, Hughes2025Trapped-ion, Ransford2025Helios] have brought this promise closer to reality, with near-term devices demonstrating potential for advantage in tasks that do not require full fault tolerance [Arute2019Quantum, Zhong2020Quantum, Kim2023Evidence, Robledo-Moreno2025Chemistry]. However, because these devices are error-prone and often remotely accessed, they raise critical questions: How can a client ensure the privacy of a delegated quantum computation while also verifying its result, especially when the output is classically intractable to compute?
This challenge is particularly acute for algorithms based on expectation value estimation, which underpin many near-term quantum applications such as quantum simulation [McArdle2020Quantum, cerezo2021variational] and machine learning [cerezo2021variational, Larocca2025Barren, Mhiri2025A, Cerezo2025Does]. For these applications, establishing a rigorous method to verify the returned estimates is a prerequisite for achieving reliable quantum utility. Yet, even in the case of honest but noisy devices, a fundamental challenge remains: is it possible for a purely classical client to efficiently and robustly verify the outcome of a task delegated to a remote server? While particular schemes have been developed for specific problems [AaronsonPeaked2024], no general practical solution currently exists for the broad class of near-term algorithms mentioned above.
(a) Round-wise verification
(b) End-to-end verification
Figure 1: (a) Schematic illustration of round-wise verification, in which the Server sends all samples directly to the Client, who then computes the empirical mean classically. In this setting, the security of each round is guaranteed to be inverse-polynomial in the number of rounds under conventional security analysis, implying that the entire protocol has at best inverse-polynomial security. (b) Schematic illustration of end-to-end verification, in which the Server returns a final empirical mean obtained via quantum averaging. In this case, exponential security can be achieved within the conventional security analysis; however, quantum averaging incurs substantial space overhead, which would be prohibitive for near-term quantum hardware. Note that in both figures, the black arrows represent quantum communication to send single qubits, and the red arrows represent the classical communication to return the final outcome from the server.
Recently, a landmark experiment [MK2025Verifiable] proposed a heuristic method where one quantum computer simulates the results of another. While this represents a major step forward, the approach is not mathematically rigorous. If two devices happen to share the same noise profile, they could produce identical results in a regime beyond classical simulability while being equally incorrect. Conversely, if the devices provide different results due to differing noise effects, there is no rigorous way to determine which result is correct.
To resolve this, researchers over the last two decades have established rigorous, scalable verification techniques based on cryptographic obfuscation. These protocols, generally referred to as Verifiable Blind Quantum Computing (VBQC) [gheorghiu2019a] are formulated as an interactive proof system. A verifier with minimal quantum capabilities, specifically single-qubit preparation, can rigorously prove the correctness of a computation performed by a potentially untrusted and noisy server. By randomly inserting traps while keeping the entire computation process obfuscated, the absence of deviations in the trap computations guarantees the correctness of the computation [fitzsimons2017unconditionally]. Such techniques require the verifier to be quantumly linked to the server, a capability demonstrated in both off-chip and on-chip settings [Gustiani2025On-chip, Polacchi2025Experimental, drmota2024verifiable]. This confirms the practicality of these schemes and their alignment with the modular scalable roadmaps of the major hardware platforms.
Nevertheless, a critical subtlety remains. To obtain exponential confidence in the correctness of the result, one must amplify the security bound. If the goal is the correctness of the entire quantum output, the only known solution is full fault-tolerant computing [fitzsimons2017unconditionally]. Fortunately for decision problems (BQP), that is the most common usage of quantum computers, one can instead blindly interleave computation rounds and test rounds, containing only traps, and apply a simple post-processing, majority voting, to achieve exponential security [leichtle2021verifying] (see Protocol 3 in Appendix B). Such a solution is then efficient as the security error is exponentially suppressed as the number of computation rounds and test rounds grows, and is unaffected by the size of the computation.
In the case of the observable estimation algorithms, the desired post-processing is instead the classical computation of the empirical average. However, this simple change breaks the existing verification framework. If the server sends all samples directly to the verifier to classically compute the average as depicted in Fig. 1(a), there is no guarantee that individual samples were correctly computed; unlike decision problems, sampling problems lack simple classical error-detection schemes. This results in a security guarantee that is at best inverse polynomial in the number of rounds. To recover exponential security guarantees, the server should perform the empirical mean quantumly by processing all samples in parallel as depicted in Fig. 1(b). Then, the space complexity of the protocol increases significantly, making the approach impractical for current hardware. This represents a fundamental gap that prevents existing verification schemes from being applicable directly to the most important class of problems targeting near-term quantum utility.
To address these limitations, we first introduce the Secure Delegated Observable Estimation (SDOE) resource. This is a new ideal resource within the abstract cryptography (AC) framework [Maurer2011abstract] that formalises a trustworthy estimation process. SDOE is specifically designed to handle the statistical nature of near-term tasks, ensuring the client obtains an estimate of an observable’s expectation value within a specified bias, $\epsilon$, or detects a deviation and aborts, all while preserving the confidentiality of the observable and the result.
We then present the Verifiable Blind Observable Estimation (VBOE) protocol, which efficiently constructs the SDOE resource. VBOE resolves the aforementioned gap by enabling a security error that is negligible in the number of rounds without requiring any extra qubit overhead. Unlike the problematic trade-offs of previous approaches, VBOE allows for the sequential collection of samples, avoiding the space blowup of parallel computation, while using a novel verification analysis to ensure the integrity of the empirical average. Our main result is a rigorous proof that VBOE constructs SDOE with composable security, providing the first efficient framework for verifying observable estimation tasks on untrusted devices.
As a consequence, our protocol directly addresses the growing demand for verifiable quantum advantage in the NISQ era. Its absence of overhead provides a path to polynomial-time verification with exponentially small soundness error for tasks performed on current state-of-the-art quantum hardware [Babbush2025The]. This bridges the gap between foundational cryptographic theory and the heuristic methods recently used in milestone experimental demonstrations of quantum advantage [Abanin2025Observation, MK2025Verifiable].
@@ II Verifying observable estimation
We define the Secure Delegated Observable Estimation (SDOE) Resource 1 and show that it can be constructed by the proposed Verifiable Blind Observable Estimation (VBOE) Protocol 1 to negligible error within the AC framework.
@@ II.1 Observable estimation problems
Using Hoeffding’s bound (Lemma 1), one can assess the performance of such a procedure:
$\begin{split}\operatorname{Pr}\left[|\mu-\operatorname{Tr}\left[\rho O\right]|\geq\epsilon\right]\leq 2\exp\left(-\frac{\epsilon^{2}}{2}N_{c}\right).\end{split}$ (2)
It states that for fixed $\epsilon$, the probability of the estimator being further away than $\epsilon$ from the true value $\operatorname{Tr}\left[\rho O\right]$ is negligible in $N_{c}$, the number of collected samples. This motivates the following definition:
@@
Given an observable $O$, a reference state $\rho$, a protocol $(\epsilon,\delta)$-estimates $\operatorname{Tr}\left[\rho O\right]$ if the protocol outputs an estimate $o$ that satisfies
$\operatorname{Pr}\left[|o-\operatorname{Tr}\left[\rho O\right]|\geq\epsilon\right]\leq\delta.$ (3)
@@ II.2 Secure delegated observable estimation (SDOE)
To formalise security for observable estimation problems, we define an ideal resource that has perfect blindness and always returns an estimate of $\operatorname{Tr}\left[\rho O\right]$ within bias $\epsilon$ in the form of Resource 1.
Resource  1 Secure Delegated Observable Estimation (SDOE)
Public information: $\mathfrak{C}$ a computation class; $N_{c},N_{t}\in\mathbb{N}$; a security parameter $\omega>0$; and the allowed bias $\epsilon>0$.
Client’s interface: The target computation $\mathsf{C}\in\mathfrak{C}$ to produce the single qubit state $\rho$ to be measured by $O=X$.
@@
    If $e=1$, it computes $s$ using the transmitted state $\sigma$ and $F$.
@@
    If $s=\mathsf{Abort}$ it forwards $o=\mathsf{Abort}$ to the Client.
@@
    If $|s-\operatorname{Tr}\left[\rho O\right]|\geq\epsilon$ it sets $o=\mathsf{Abort}$ and forwards it to the Client.
@@
Here, $e\in\{0,1\}$ is a flag controlling whether the Server’s interface filter is activated or not. Whenever $e=0$, the ideal resource samples the estimator of $\operatorname{Tr}\left[\rho O\right]$ and returns its value if it is within $\epsilon$ of the true expectation value. When the Server asks for full access ($e=1$), the Server receives at most the permitted leakage, i.e. essentially $\mathfrak{C}$ corresponding to all the observable estimation problems that the resource can handle. It also recieves the parameters $N_{c}$, $N_{t},w$ and $\epsilon$. The Server is then allowed to send a deviation to be applied by the Resource. It takes the form of a quantum state $\sigma$ and a classical list of quantum and classical instructions that produce either a real number or $\mathsf{Abort}$. If the produced scalar is within $\epsilon$ of $\operatorname{Tr}\left[\rho O\right]$, then the resource sends the scalar to the Client. Otherwise, it sends $\mathsf{Abort}$.
This definition corresponds to the intuitive notion of secure delegated observable estimation: a malicious Server can only learn the class of observable estimation problems that the resource can tackle, and is only able to influence the result so long as it remains within $\epsilon$ of the true expectation value. It also inevitably has the ability to force the protocol to $\mathsf{Abort}$. The pictorial representation of the SDOE resource is provided in Fig. 2.
@@ II.3 Verifiable blind observable estimation (VBOE)
Considering the SDOE resource defined above, we propose the VBOE protocol to concretely implement this functionality. It is based on the same sequential execution of test and computation rounds as in the RVBQC (See appendix, Protocol 3[leichtle2021verifying], with post-processing and acceptance criteria tailored to observable estimation problems. Here we assume that $\mathfrak{C}$ is a computation class that corresponds to all measurement patterns that can be executed on a graph $G=(V,E)$ with a given flow $f$.
Protocol  1 Verifiable Blind Observable Estimation (VBOE)
Inputs from Client: The target computation $\mathsf{C}\in\mathfrak{C}$ that produces $\rho$ and allows to measure $\operatorname{Tr}\left[\rho O\right]$ with $O=X$.
@@
    For $i\in\mathrm{S}_{\mathsf{T}}$, the Client constructs a test round following the same procedure as for the test rounds of the RVBQC (Step 2 of Protocol 3).
@@
    Each round, computation or test, is then delegated to the Server using UBQC (Protocol 2).
@@
    Upon receiving and decoding the result of the computation round $i\in\mathrm{S}_{\mathsf{C}}$, the Client assigns the result to $\tilde{y}_{i}\in\{-1,1\}$.
@@
    If less than $\omega N_{t}$ test rounds failed, it sets $\tilde{o}=\frac{1}{N_{c}}\sum_{i\in\mathrm{S}_{\mathsf{C}}}\tilde{y}_{i}$ as the result, otherwise it sets it to $\tilde{o}=\mathsf{Abort}$.
The main modification in VBOE from RVBQC is the post-processing of computation rounds. RVBQC applies classical majority vote over multiple repeated runs to amplify the probability of choosing a correct outcome, while VBOE computes the empirical average over $N_{c}$ outcomes $\{\tilde{y}_{i}\}_{i\in\mathrm{S}_{\mathsf{C}}}$. To ensure that the returned value stays within the allowed bound $\epsilon$, the threshold $\omega$ needs to be set at the appropriate value. More precisely, we will see in the security proof that it is set in a way that the concentration of probabilities ensures honest executions are always accepted, while deviations that could generate a result too far away from the true value are rejected. The schematic illustration of the VBOE protocol is depicted in Fig. 3.
@@ II.4 Concrete construction of the SDOE resource
Let $\mathfrak{C}$ be a class of observable estimation problems that can be estimated using an MBQC pattern on a fixed graph $G$ with a given flow $f$ and chromatic number $K$. Let $N_{c},N_{t},\in\mathbb{N}$, let $\epsilon,\omega$ be constants such that $0\leq K\omega<\epsilon$. Then, the VBOE protocol (Protocol 1), with $N_{c}$ computation rounds and $N_{t}$ test rounds, $\delta$-constructs the SDOE resource. For a constant ratio $N_{c}/N_{t}$, $\delta$ is negligible in $N_{c}$.
Following abstract cryptography, to prove this theorem, we need to upper bound the distinguishing advantage between the VBOE protocol and the SDOE resource in the honest (Correctness proof) and malicious (Security proof) settings (see Definition 4 in Appendix B).
@@
The proof of correctness relies on the composability of the UBQC protocol. As apparent in Protocol 1, each round is delegated to the server using UBQC (Protocol 2 in Appendix B). Because UBQC perfectly constructs the Blind Delegated Quantum Computation (BDQC) resource (Resource 2 in Appendix B), we can instead perform the proof of correctness using a hybrid protocol where each instantiation of UBQC is replaced by a call to BDQC.
As a result, each outcome $y_{i}$ for the computation rounds $i\in\mathrm{S}_{\mathsf{C}}$ is processed through $y_{i}=2Y_{i}-1$, where $Y_{i}$ is sampled from the Bernoulli distribution $\mathcal{B}(p)$ with probability $p=\frac{1+\operatorname{Tr}\left[\rho O\right]}{2}$. This ensures that the produced empirical average $\tilde{\mu}=\frac{1}{N_{c}}\sum_{i\in\mathrm{S}_{\mathsf{C}}}\tilde{y}_{i}$ is obtained from the same probability distribution as the one used to define the ideal resource. Hence, whenever the ideal resource and the protocol both output the estimated value or both output $\mathsf{Abort}$, their outputs coincide. Consequently, the only distinguishing advantage stems from the two setups having different $\mathsf{Abort}$ probabilities.
Indeed, because the test rounds in the protocol are executed perfectly, they consistently give the correct outcome, and the protocol never aborts. For the ideal resource, this is not the case. Whenever the estimator is further away than $\epsilon$ from $\operatorname{Tr}\left[\rho O\right]$, the ideal resource returns $\mathsf{Abort}$. The probability of such an event happening is upper bounded, using Hoeffding’s bound, by $2\exp\left(-\frac{\epsilon^{2}}{2}N_{c}\right)$. We can thus conclude that, for $\epsilon$ fixed, the distinguishing advantage in the honest case is a negligible function of $N_{c}$. ∎
@@
The security relies heavily on the composability of UBQC. We start by constructing a simulator that we attach to the Server’s interface of the ideal resource. Its purpose is to generate plausible transcripts and help the ideal resource in returning $o\in\mathbb{R}\cup\{\mathsf{Abort}\}$ to the Client’s interface so that any distinguisher will be unable to tell apart this situation from running the Client’s part of the concrete Protocol 1. This simulator is easily constructed from the simulator designed to prove the security of UBQC.
First, it sets $e=1$. It then prepares EPR pairs for each qubit that the Server is supposed to receive in Protocol 1. It sends all half EPR pairs to the Server, instructs random measurement angles, and retrieves alleged measurement outcomes. It then forwards the second half of the EPR pairs, the chosen angles and received bits to the ideal resource. It also samples at random indices within $N_{c}+N_{t}]$ to define the sets $\mathrm{S}_{\mathsf{C}}$ and $\mathrm{S}_{\mathsf{T}}$. It decides which type of test round is associated with each $i\in\mathrm{S}_{\mathsf{T}}$ and passes this to the ideal resource.
Following the security proof of UBQC, the information passed per round, together with the nature of the round, either computation or test, is sufficient for the ideal resource to generate per-round measurement results following the same, possibly deviated, probability distributions as the ones obtained when running Protocol 1. From the outcomes of computation rounds, the ideal resource is then instructed to compute the empirical average $o$ and from the test rounds to check that no more than $\omega N_{t}$ failed, in which case it sets $o=\mathsf{Abort}$. Using the computed value $o$, the ideal resource then performs the steps 3, 4, and 5 mentioned in its definition (Resource 1) and aimed at ensuring that the returned $o$ is within $\epsilon$ of $\operatorname{Tr}\left[\rho O\right]$.
To proceed further with the proof, we note that the perfect blindness of UBQC ensures that the value of $s$ computed by using the quantum state $\sigma$ provided by the simulator, i.e. the half EPR pairs, and the classical instructions follows the same distribution as $\tilde{o}$ in Protocol 1. The only difference arrives later when the additional check $|s-\operatorname{Tr}\left[\rho O\right]|\geq\epsilon$ is performed by the resource and possibly rejects when the protocol would have accepted.
As a result, the distinguishing advantage for telling apart the concrete protocol from the ideal resource is bounded by the total variation distance between the probability distributions of $s$ and $o$, or equivalently between $o$ and $\tilde{o}$. Because, conditioned on $o,\tilde{o}\in\mathbb{R}$ or $o,\tilde{o}=\mathsf{Abort}$, the distributions of $o$ and $\tilde{o}$ are the same, the total variation distance reduces to:
$\begin{split}&|\operatorname{Pr}[o=\mathsf{Abort}]-\operatorname{Pr}[\tilde{o}=\mathsf{Abort}]|\\ &\quad\quad=\operatorname{Pr}[\tilde{o}\neq\mathsf{Abort}\wedge|\tilde{o}-\operatorname{Tr}\left[\rho O\right]|\geq\epsilon].\end{split}$ (4)
@@
    we upper bound the probability that the computation rounds provide an empirical average that is more than $\gamma_{1}$ away from $\operatorname{Tr}\left[\rho O\right]$, with $\gamma_{1}>0$;
@@
    given that $\tilde{o}\neq\mathsf{Abort}$, we then upper-bound the probability that a large number of computation rounds, say $(K\omega+\gamma_{2})N_{c}$ for $\gamma_{2}>0$, have been attacked by the server;
@@
    we recognise that attacking a fraction $\phi$ of computation rounds yields a deviated empirical average that is not away by more than $2\phi$ from the non-deviated one, due to the binary nature of the observable $O$, taking either 1-1 or $1$ for each round;
@@
    we then notice that the probabilities in steps 1 and 2 above are negligible functions of $N_{c}$ and $N_{t}$. Thus, if we set $\gamma_{1}+2(K\omega+\gamma_{2})<\epsilon$, we have $\operatorname{Pr}\left[\tilde{o}\neq\mathsf{Abort}\wedge|\tilde{o}-\operatorname{Tr}\left[\rho O\right]|\geq\epsilon\right]$ negligible in $N_{c}$ and $N_{t}$, as it is upper bounded by the sum of the probabilities of step 1 and 2 as a result of the union bound.
@@
First, it is straightforward to see that the probability at step 1 is upper-bounded by $2\exp\left(-\frac{\gamma_{1}^{2}}{2}N_{c}\right)$ using Hoeffding’s inequality.
@@
The probability of step 2 is upper-bounded in the following way. First, because UBQC reduces the attack by the Server to a convex combination of Pauli deviations before the measurements, we can group the deviation strategies by $m$, the number of attacked rounds. Now, because the location of test rounds and computation rounds are random, the number of affected computation rounds $Z$ follows a hypergeometric distribution with parameters $N_{c}+N_{t}$ for the total number of items, $m$ for the number of marked items and $N_{c}$ for the number of samples. Similarly, the number of affected test rounds, denoted by $X$, follows a hypergeometric distribution with the roles $N_{c}$ and $N_{t}$ swapped.
@@
Let $Y$ be the number of failed test rounds. The idea is now to upper-bound the probability $\operatorname{Pr}[Z\geq(K\omega+\gamma_{2})N_{c},Y\leq\omega N_{t}]$ using the following fact that